solve the system by elimination. is it inconsistant or dependent?
3x + 2y = 5
-12x - 8y = -17

Answer by mananth(16949)

3x+2y=5 .............1
-12x-8y =-17 .............2

multiply (1)by 4
Multiply (2) by 1
12x+8y =20
-12x-8y =-17

Add the two equations

0x+0y =3
The system has no solution.
Hence inconsistent.